# ThinkEdit: Interpretable Weight Editing to Mitigate Overly Short Thinking in Reasoning Models

### v1

@@ 3.3 Effects of Reasoning-Length Direction
We think the steering representation approach offers a more structured and flexible method for controlling the inference budget compared with prior work [15]. In [15], the authors enforce budget constraints by prematurely stopping the CoT (often causing incomplete reasoning) or appending "Wait" to artificially prolong it (risking misalignment with the model’s natural CoT); In contrast, our approach of steering representations allow for a more coherent way to modulate reasoning length (see Section A.2). By directly manipulating the model’s internal representations, one can more effectively balance the computational cost and performance.
@@ 4.4 Performance of Reasoning Models after ThinkEdit
Table 1: Overall accuracy (%) of each model before and after attention-weight editing.
@@
Finally, Table 3 reports the average reasoning length among the top 5%, 10%, and 20% shortest responses. We observe that ThinkEdit consistently increases the reasoning length in these short-answer scenarios, suggesting that the intervention discourages excessively short reasoning, and therefore leads to higher accuracy as shown in Table 2.

### v2

@@ 3.3 Effects of Reasoning-Length Direction
Recent work [15] proposed an interesting approach to enforce budget constraints by stopping the CoT or appending "Wait" to prolong it. However, stopping the CoT prematurely may cause incomplete reasoning and appending "Wait" may risk misalignment with the model’s natural CoT. Alternatively, steering representations may allow for a more coherent way to modulate reasoning length (see Section A.2) – by directly manipulating the model’s internal representations, one can more effectively balance the computational cost and performance.
@@ 4.4 Performance of Reasoning Models after ThinkEdit
Table 1: Overall accuracy (%) of each model before and after applying ThinkEdit.
@@
Finally, Table 3 reports the average reasoning length among the top 5%, 10%, and 20% shortest responses. We observe that ThinkEdit consistently increases the reasoning length in these short-answer scenarios, suggesting that the intervention discourages excessively short reasoning, and therefore leads to higher accuracy as shown in Table 2. Interestingly, despite the local increase, the average reasoning length across all responses changes by +3.5% for deepseek-qwen-14b, +3.6% for deepseek-llama3-8b, and -6.7% for deepseek-qwen-1.5b. As a result, the overall net change in reasoning length across all models is a slight decrease of 0.5%. This suggests that ThinkEdit selectively extends short reasoning traces without unnecessarily prolonging already long ones.